# Visibly Fair Mechanisms

### v1

@@ A.4 Proof of Theorem 4
If part a. A visibly fair mechanism satisfying weak availability is strategy-proof if it satisfies expressiveness. We prove the contrapositive statement: If a visibly fair mechanism satisfying weak availability is not strategy-poof it does not satisfy expressiveness.
@@
    By assumption, there exists a truthful message $m_{i}$ and a message $\hat{m}_{i}\in M_{i}$ such that $\psi(\hat{m}_{i},m_{-i})_{i}\succ_{i}\psi(m)_{i}$.
@@
    By weak availability $\psi(\hat{m}_{i},m_{-i})_{i}$ is available under $\psi(m)$.
@@
    The information $\psi(\hat{m}_{i},m_{-i})_{i}\succ_{m_{i}}\psi(m)_{i}$ would lead to a violation of visible fairness. Thus, we have found a violation of expressiveness.
If part b. A visibly fair mechanism satisfying expressiveness is strategy-proof if it satisfies weak availability. We prove the contrapositive statement: If a visibly fair mechanism satisfying expressiveness is not strategy-poof it does not satisfy weak availability.
@@
    By assumption, there exists a truthful message $m_{i}$ and a message $\hat{m}_{i}\in M_{i}$ such that $\psi(\hat{m}_{i},m_{-i})_{i}\succ_{i}\psi(m)_{i}$.
@@
    By expressiveness the information $\psi(\hat{m}_{i},m_{-i})_{i}\succ_{m_{i}}\psi(m)_{i}$ is available under $\psi(m)$.
@@
    The availability of $\psi(\hat{m}_{i},m_{-i})_{i}$ under $\psi(m)$ would lead to a violation of visible fairness. Thus, we have found a violation of weak availability.
Only if part a. A visibly fair mechanism satisfying weak availability also satisfies expressiveness if it is strategy-proof. We prove the contrapositive statement: If a visibly fair mechanism satisfying weak availability does not satisfy expressiveness then it is not strategy-proof.
@@
    By assumption there exist messages $m_{i}$ and $\hat{m}_{i}$ s.t. neither $\psi(\hat{m}_{i},m_{-i})_{i}\succ_{m_{i}}\psi(m)_{i}$ nor $\psi(m)_{i}\succ_{m_{i}}\psi(\hat{m}_{i},m_{-i})_{i}$
@@
Only if part b. A visibly fair mechanism satisfying expressiveness also satisfies weak availability if it is strategy-proof. We prove the contrapositive statement: If a visibly fair mechanism satisfying expressiveness does not satisfy weak availability then it is not strategy-proof.
@@
    As $m_{i}\in M_{i}$ that is truthful for a preference $\succsim_{i}\in\mathcal{Q}_{i}$ it directly follows that $m^{\prime}_{i}\in M_{i}\setminus\{m_{i}\}$ is a successful manipulation as $\psi(m^{\prime}_{i},m_{-i})_{i}\succ_{i}\psi(m)_{i}$.
@@
A mechanism is strategy-proof if and only if it satisfies consistency.
@@
If part a. A mechanism is strategy-proof if it satisfies consistency. We prove the contrapositive statement: If a mechanism is not strategy-poof it does not satisfy expressiveness.
By assumption, there exists a truthful message $m_{i}$ and a message $\hat{m}_{i}\in M_{i}$ such that $\psi(\hat{m}_{i},m_{-i})_{i}\succ_{i}\psi(m)_{i}$. As $\psi(\hat{m}_{i},m_{-i})_{i}$ is strictly preferred we have $\psi(\hat{m}_{i},m_{-i})_{i}\neq\psi(m)_{i}$. As $m_{i}$ is truthful we cannot have $\psi(m)\succ_{m_{i}}\psi(\hat{m}_{i},m_{-i})$, and therefore $\psi(\hat{m}_{i},m_{-i})\in G(\{\psi(m)\}\cup\{\psi(\hat{m}_{i},m_{-i})\},m_{i})$, i.e., a violation of consistency.
Only if part a. A mechanism satisfies consistency if it is strategy-proof. We prove the contrapositive statement: If a mechanism does not satisfy consistency then it is not strategy-proof.
By assumption, there exists a message $m_{i}$ and a message $\hat{m}_{i}\in M_{i}$ such that $\psi(\hat{m}_{i},m_{-i})\in G(\{\psi(m)\}\cup\{\psi(\hat{m}_{i},m_{-i})\}$. Consider any preference $\succsim^{*}_{i}\in Q_{i}$ s.t. $\psi(\hat{m}_{i},m_{-i})\succ^{*}_{i}\psi(m)$ and $s\succ^{*}_{i}s^{\prime}$ whenever $s\succ_{m_{i}}s^{\prime}$. By construction, $m_{i}$ is a truthful message for preference $\succsim^{*}_{i}\in Q_{i}$. Moreover, there exists another message $\hat{m}_{i}\in M_{i}$ such that $\psi(\hat{m}_{i},m_{-i})_{i}\succ_{i}\psi(m)_{i}$, i.e., a violation of strategy-proofness.
@@ A.5 Proof of Theorem 6
First, it should be clear that the Modular Priority Mechanism is a partitioned priority mechanism: each agent, following the priority order, is associated with a zonal message space, and is matched to the most-preferred state from a zone that still has states with spare capacity.
Next, we show that, for any given set of type-specific modular upper-bound systems and exogenous rankings $(\blacktriangleright_{i})_{i\in I}$, the Modular Priority Mechanism, represented by the function $\psi$,is strategy-proof. Notice first that, by construction, an officer $i$ cannot, by submitting a different message, change the assignment of any officer $j<i$ and, as a consequence, the zones that are feasible for her. Since the zone from which $i$’s assignment will be drawn from depends on the assignment of officers with higher priority and the ranking $\blacktriangleright_{i}$, both unaffected by $m_{i}$.
Notice, therefore, that since for any $m$ and $m_{i}^{\prime}$, $\psi(m)$ and $\psi(m_{i}^{\prime},m_{-i})$ are comparable under both $m_{i}$ and $m_{i}^{\prime}$ (since they must all be in the same zone).This implies that, for any officer $i$ and true preference $\succsim_{i}\in\mathcal{Q}_{i}$, if $m_{i}$ is a truthful message, then $m_{i}$ completely ranks all states in a zone exactly like $\succsim_{i}$.
All of this implies that there is no preference $\succsim_{i}\in\mathcal{Q}_{i}$ with truthful message $m_{i}\in M_{i}$ and $m^{\prime}_{i}\in M_{i}\setminus\{m_{i}\}$ for which $\psi(m_{i}^{\prime},m_{-i})\neq\psi(m)$ and $\psi(m_{i}^{\prime},m_{-i})\succ_{i}\psi(m)$. Therefore, $\psi$ trivially satisfies expressiveness and availability.
Finally, we need to show that the Modular Priority Mechanism respects modular upper-bounds. That is, the final allocation $\psi(m)=a$ satisfies, for every type $t\in T$ and every upper-bound $(S_{h}^{t},k_{h}^{t})\in H^{t}$:
|$\left|\left\{i\in I:a_{i}\in S_{h}^{t},\ t_{i}=t\right\}\right|\leq k_{h}^{t}.$
We prove this by induction on the steps $k=1,2,\ldots,n$ of the mechanism. In order to facilitate notation and comprehension, we will denote by $a^{t}$ the “tentative” allocation $a$ by the end of step $t$ of the mechanism.
@@
At initialization, no officers have been assigned. That is, $a^{0}_{i}=\varnothing$ for all $i\in I$. Therefore, for any type $t\in T$ and upper-bound $(S_{h}^{t},k_{h}^{t})\in H^{t}$, we have:
|$\left|\left\{i\in I:a^{0}_{i}\in S_{h}^{t},\ t_{i}=t\right\}\right|=0\leq k_{h}^{t}.$
@@
Assume that at step $k-1$, the current assignment $a^{k-1}$ respects all modular upper-bounds; that is, for every type $t\in T$ and upper-bound $(S_{h}^{t},k_{h}^{t})\in H^{t}$:
$N_{h}^{k-1}=\left|\left\{i\in I:a_{i}^{k-1}\in S_{h}^{t},\ t_{i}=t\right\}\right|\leq k_{h}^{t}.$
@@
Consider officer $i_{k}$ of type $t_{i_{k}}$. There are two cases:
Case 1: Officer $i_{k}$ is assigned to a state $s^{\ast}$:
According to the mechanism, $i_{k}$ is assigned to a state $s^{\ast}$ within a zone $z^{*}$ where the upper-bounds for type $t_{i_{k}}$ involving states in that zone are not yet binding before $i_{k}$’s assignment—this is tracked by the flag $B_{z^{*}}^{t_{i_{k}}}=\text{False}$.
That is, for every upper-bound involving type $t_{i_{k}}$ and upper-bound $(S_{h}^{t},k_{h}^{t})\in H^{t}$ for which $s^{\ast}\in S_{h}^{t}$:
|$\left|\left\{i\in I:a_{i}^{k-1}\in S_{h}^{t},\ t_{i}=t\right\}\right|<k_{h}^{t}.$
@@
Case 2: Officer $i_{k}$ remains unmatched:
If no feasible zone is found for $i_{k}$, they remain unmatched; that is, $a^{k}_{i_{k}}=\varnothing$ and $a^{k}_{i}=a^{k-1}_{i}$ for all $i\neq i_{k}$. In this case, $a^{k-1}_{i}$ respects upper-bounds, so does $a^{k}_{i}$.
@@ Conclusion:
|$\left|\left\{i\in I:a_{i}^{n}\in S_{h}^{t},\ t_{i}=t\right\}\right|\leq k_{h}^{t},$
for every type $t\in T$ and every upper-bound $(S_{h}^{t},k_{h}^{t})\in H^{t}$.
@@ A.6 Proof of Theorem 7
Consider a problem with three officers $I=\{i_{1},i_{2},i_{3}\}$ (with priority $\pi(i_{1})<\pi(i_{2})<\pi(i_{3})$) and two states $S=\{s_{1},s_{2}\}$ with capacity $q_{s_{1}}=q_{s_{2}}=2$. Assume that all officers are of the same type $t$ and that the modular upper–bound for type $t$ is ($(\{s_{1}\},1)$; that is, at most one officer of type $t$ may be assigned to $s_{1}$.
@@
Next, note that no mechanism that elicits preferences of more than one officer between $s_{1}$ and $s_{2}$ and is visibly fair respects modular upper-bounds. To see that, assume that two officers express that $s_{1}\succ_{i}s_{2}$. Visible fairness requires that $s_{1}$ does not have and empty slot. But this would violate modular upper-bounds.
@@
Case 2: $(X_{1},X_{2},X_{3})=(\text{N},\text{N},\text{Y})$. Suppose that $i_{3}$ reports $s_{2}\succ_{i_{3}}s_{1}$. There are two cases to consider: (i) one of $i_{1}$,$i_{2}$ is matched to $s_{1}$, or (ii) both $i_{1}$ and $i_{2}$ are matched to $s_{2}$. In the first case, let without loss of generality $i_{1}$ be matched to $s_{2}$ and $i_{2}$ be matched to $s_{1}$ and. Then, if $i_{1}$ prefers $s_{1}$ and $i_{2}$ prefers $s_{2}$, swapping their assignments would Pareto dominate the outcome produced by the mechanism while still respecting modular upper-bounds.
@@
In every case, the following occurs: either the outcome produced by the mechanism (which is constrained to be upper–bound respecting) violates the modular upper–bound (by assigning more than one officer to $s_{1}$), or there exists another allocation that both respects the upper–bounds and yields a strict Pareto improvement with respect to the officers’ true preferences. Therefore, no mechanism exists that is simultaneously visibly fair and is constrained Pareto efficient.
@@ A.7 Proof of Theorem 8
We show by induction on the assignment order that for any type $t\in T$ and any modular upper-bound $(S_{h}^{t},k_{h}^{t})$, the number of type-$t$ officers assigned to states in $S_{h}^{t}$ never exceeds $k_{h}^{t}$.
Base Case. Before any assignment (step 0), no officer is assigned and the count is $0\leq k_{h}^{t}$.
Inductive Step. Suppose that after assigning officers $i_{1},\dots,i_{k-1}$, the upper-bound $(S_{h}^{t},k_{h}^{t})$ is not violated. When officer $i_{k}$ (of type $t_{i_{k}}$) is considered, if $t_{i_{k}}\neq t$ the claim is unaffected. If $t_{i_{k}}=t$ and $i_{k}$ is assigned a state $s$ with $s\in S_{h}^{t}$, then by construction $s$ belongs to zone $z_{1}$ for $i_{k}$ and the current count of type-$t$ officers in $S_{h}^{t}$ is strictly less than $k_{h}^{t}$. If after assignment the count reaches $k_{h}^{t}$, the mechanism updates $B^{t}$ by adding $(S_{h}^{t},k_{h}^{t})$; hence, any subsequent officer of type $t$ will have $S_{h}^{t}$ excluded from her zone $z_{1}$. Thus, by induction, no upper-bound is ever violated.
@@
    Binding Upper-Bound: The other reason for $a^{\prime}_{i^{*}}$ not to be in $i^{*}$’s zone $z_{1}$ under $a$ is if $a^{\prime}_{i^{*}}$ belonged to a set $S_{h}^{t_{i^{*}}}$ for which the corresponding upper-bound $(S_{h}^{t_{i^{*}}},k_{h}^{t_{i^{*}}})$ was already binding at the time of $i^{*}$’s assignment. However, if $a^{\prime}_{i^{*}}$ were in such a set and $a^{\prime}$ respects the upper-bound, then $a^{\prime}_{i^{*}}$ could not be assigned to $i^{*}$ without causing the total number of type $t_{i^{*}}$ officers in $S_{h}^{t_{i^{*}}}$ to exceed $k_{h}^{t_{i^{*}}}$.

### v2

@@ A.4 Proof of Theorem 4
If part a. A visibly fair mechanism satisfying weak availability is strategy-proof if it satisfies expressiveness. We prove the contrapositive statement: If a visibly fair mechanism satisfying weak availability is not strategy-proof it does not satisfy expressiveness.
@@
    By assumption (violation of strategy-proofness), there exists a truthful message $m_{i}$ and a message $\hat{m}_{i}\in M_{i}$ such that $\psi(\hat{m}_{i},m_{-i})_{i}\succ_{i}\psi(m)_{i}$.
@@
    By weak availability $\psi(\hat{m}_{i},m_{-i})_{i}$ is available under \$m\$.
@@
    The expressed information $\psi(\hat{m}_{i},m_{-i})_{i}\succ_{m_{i}}\psi(m)_{i}$, together with $\psi(\hat{m}_{i},m_{-i})_{i}$ being available under $m$, would lead to a violation of visible fairness. Thus, we have found a violation of expressiveness.
If part b. A visibly fair mechanism satisfying expressiveness is strategy-proof if it satisfies weak availability. We prove the contrapositive statement: If a visibly fair mechanism satisfying expressiveness is not strategy-proof it does not satisfy weak availability.
@@
    By assumption (violation of strategy-proofness), there exists a truthful message $m_{i}$ and a message $\hat{m}_{i}\in M_{i}$ such that $\psi(\hat{m}_{i},m_{-i})_{i}\succ_{i}\psi(m)_{i}$.
@@
    By expressiveness the information $\psi(\hat{m}_{i},m_{-i})_{i}\succ_{m_{i}}\psi(m)_{i}$ is available under $m$.
@@
    The availability of $\psi(\hat{m}_{i},m_{-i})_{i}$ under $m$, together with the expressed information $\psi(\hat{m}_{i},m_{-i})_{i}\succ_{m_{i}}\psi(m)_{i}$, would lead to a violation of visible fairness. Thus, we have found a violation of weak availability.
Only if part a. A visibly fair mechanism satisfies expressiveness if it is strategy-proof. We prove the contrapositive statement: If a visibly fair mechanism does not satisfy expressiveness then it is not strategy-proof.
@@
    Note that for the same state expressiveness is always satisfied. As the mechanism does not satisfy expressiveness, there exist officer $i$ and messages $m_{i}$ and $\hat{m}_{i}$ such that $\psi(m)_{i}\neq\psi(\hat{m}_{i},m_{-i})_{i}$, but neither $\psi(\hat{m}_{i},m_{-i})_{i}\succ_{m_{i}}\psi(m)_{i}$ nor $\psi(m)_{i}\succ_{m_{i}}\psi(\hat{m}_{i},m_{-i})_{i}$
@@
Only if part b.
A visibly fair mechanism satisfies weak availability if it is strategy-proof. We prove the contrapositive statement: If a visibly fair mechanism does not satisfy weak availability then it is not strategy-proof.
@@
    As the mechanism does not satisfy weak availability, there exists $m_{i}\in M_{i}$ that is truthful for a preference $\succsim_{i}\in\mathcal{Q}_{i}$ and a message $m^{\prime}_{i}\in M_{i}\setminus\{m_{i}\}$ s.t. $\psi(\hat{m}_{i},m_{-i})_{i}\succsim_{i}\psi(m)_{i}$; but state $\psi(\hat{m}_{i},m_{-i})_{i}$ is not available to officer $i$ under $m$.
  2. 2.
    It follows that $\psi(\hat{m}_{i},m_{-i})_{i}\neq\psi(m)_{i}$, and therefore we have a successful manipulation, as $\psi(m^{\prime}_{i},m_{-i})_{i}\succ_{i}\psi(m)_{i}$.
@@
A mechanism is strategy-proof if and only if it satisfies coherence.
@@
If. A mechanism is strategy-proof if it satisfies coherence. We prove the contrapositive statement: If a mechanism is not strategy-proof it does not satisfy coherence.
By assumption, there exists a truthful message $m_{i}$ and a message $\hat{m}_{i}\in M_{i}$ such that $\psi(\hat{m}_{i},m_{-i})_{i}\succ_{i}\psi(m)_{i}$. As $\psi(\hat{m}_{i},m_{-i})_{i}$ is strictly preferred we have $\psi(\hat{m}_{i},m_{-i})_{i}\neq\psi(m)_{i}$. As $m_{i}$ is truthful we cannot have $\psi(m)\succ_{m_{i}}\psi(\hat{m}_{i},m_{-i})$, and therefore $\psi(\hat{m}_{i},m_{-i})\in G(\{\psi(m)\}\cup\{\psi(\hat{m}_{i},m_{-i})\},m_{i})$, i.e., a violation of coherence.
Only if. A mechanism satisfies coherence if it is strategy-proof. We prove the contrapositive statement: If a mechanism does not satisfy coherence then it is not strategy-proof.
By assumption, there exists a message $m_{i}$ and a message $\hat{m}_{i}\in M_{i}$ such that $\psi(\hat{m}_{i},m_{-i})\in G(\{\psi(m)\}\cup\{\psi(\hat{m}_{i},m_{-i})\},m_{i})$. Consider any preference $\succsim^{*}_{i}\in Q_{i}$ s.t. $\psi(\hat{m}_{i},m_{-i})\succ^{*}_{i}\psi(m)$ and $s\succ^{*}_{i}s^{\prime}$ whenever $s\succ_{m_{i}}s^{\prime}$. By construction, $m_{i}$ is a truthful message for preference $\succsim^{*}_{i}\in Q_{i}$. Moreover, there exists another message $\hat{m}_{i}\in M_{i}$ such that $\psi(\hat{m}_{i},m_{-i})_{i}\succ_{i}\psi(m)_{i}$, i.e., a violation of strategy-proofness.
@@ A.5 Proof of Theorem 6
The Modular Priority Mechanism (MPM) is visibly fair, strategy-proof and respects modular upper-bounds.
@@
First, it should be clear that the Modular Priority Mechanism is a partitioned priority mechanism: each agent, following the priority order, is associated with a zonal message space, and is matched to the most-preferred state from a zone that still has states with spare capacity. Since the modular upper‑bound system $H$ satisfies sequential solvency, at every step of the MPM an officer has a state available for which no upper-bound is binding and with spare capacity.
Next, we show that, for any given set of type-specific modular upper-bound systems and exogenous rankings $(\blacktriangleright_{i})_{i\in I}$, the Modular Priority Mechanism, represented by the function $\psi$, is strategy-proof.
Notice first that, by construction, an officer $i$ cannot, by submitting a different message, change the assignment of any officer $j<i$. Since $\psi$ only produces feasible outcomes, this implies that $\psi$ satisfies availability.
Since the zone $z$ from which $i$’s assignment will be drawn from depends on the assignment of officers with higher priority and the ranking $\blacktriangleright_{i}$, both unaffected by $m_{i}$, the zone to which $i$ will be assigned during the execution of the mechanism cannot be changed by changes in her message. Let $s=\psi(m)$. Since the zone from which $i$’s outcome will be drawn will always be in $z$ regardless of the message $i$ sends, $\psi(m_{i}^{\prime},m_{-i})\in z$ for any $m_{i}^{\prime}\in M_{i}$. Since both $s$ and $\psi(m_{i}^{\prime},m_{-i})$ belong to zone $z$, it must be the case that $s$ and $\psi(m_{i}^{\prime},m_{-i})_{i}$ are comparable under $m_{i}$. Therefore, $\psi$ satisfies expressiveness, and by Corollary 2, strategy-proofness.
Finally, we need to show that the Modular Priority Mechanism respects modular upper-bounds. That is, the final allocation $\psi(m)=a$ satisfies, for every upper-bound $(\Xi_{h},S_{h},k_{h})\in H$:
$\left|\left\{i\in I:a_{i}\in S_{h},\ t_{i}\in\Xi_{t}\right\}\right|\leq k_{h}.$
We prove this by induction on the steps $k=1,2,\ldots,n$ of the mechanism. In order to facilitate notation and comprehension, we will denote by $a^{k}$ the “tentative” allocation $a$ by the end of step $k$ of the mechanism.
@@
At initialization, no officers have been assigned. That is, $a^{0}_{i}=\varnothing$ for all $i\in I$. Therefore, for any upper-bound $(\Xi_{h},S_{h},k_{h})\in H$, we have:
$\left|\left\{i\in I:a^{0}_{i}\in S_{h},\ t_{i}\in\Xi_{h}\right\}\right|=0\leq k_{h}.$
@@
Assume that at step k1k-1, the current assignment $a^{k-1}$ respects all modular upper-bounds; that is, for every upper-bound $(\Xi_{h},S_{h},k_{h})\in H$:
$N_{h}^{k-1}=\left|\left\{i\in I:a_{i}^{k-1}\in S_{h},\ t_{i}\in\Xi_{h}\right\}\right|\leq k_{h}.$
@@
Consider officer $i_{k}$ of type $t_{i_{k}}$. According to the mechanism, $i_{k}$ is assigned to a state $s^{\ast}$ within a zone $z^{t_{i}}_{\ell^{*}}$ where the upper-bounds for type $t_{i_{k}}$ involving states in that zone are not yet binding before $i_{k}$’s assignment—this is tracked by the flag $B_{\ell^{*}}^{t_{i_{k}}}=\text{False}$.
That is, for every upper-bound $(\Xi_{h},S_{h},k_{h})\in H$ for which $s^{\ast}\in S_{h}$:
$\left|\left\{i\in I:a_{i}^{k-1}\in S_{h},\ t_{i}\in\Xi_{h}\right\}\right|<k_{h}.$
@@ Conclusion:
$\left|\left\{i\in I:a_{i}^{n}\in S_{h},\ t_{i}\in\Xi_{h}\right\}\right|\leq k_{h},$
for every every upper-bound $h\in H$.
@@ A.6 Proof of Theorem 7
Not all modular upper-bound constraints admit a (static) mechanism that is simultaneously visibly fair and constrained Pareto efficient.
@@
Consider a problem with three officers $I=\{i_{1},i_{2},i_{3}\}$ (with priority $\pi(i_{1})<\pi(i_{2})<\pi(i_{3})$) and two states $S=\{s_{1},s_{2}\}$ with capacity $q_{s_{1}}=q_{s_{2}}=2$. Assume that all officers are of the same type $t$ and that there is a single modular upper–bound $(\{t\},\{s_{1}\},1)$; that is, at most one officer of type $t$ may be assigned to $s_{1}$.
@@
Next, note that no mechanism that elicits preferences of more than one officer between $s_{1}$ and $s_{2}$ and is visibly fair respects modular upper-bounds. To see that, assume that two officers express that $s_{1}\succ_{i}s_{2}$. Visible fairness requires that $s_{1}$ does not have any empty slot. But this would violate modular upper-bounds.
@@
Case 2: $(X_{1},X_{2},X_{3})=(\text{N},\text{N},\text{Y})$. Suppose that $i_{3}$ reports $s_{2}\succ_{i_{3}}s_{1}$. There are two cases to consider: (i) one of $i_{1}$$i_{2}$ is matched to $s_{1}$, or (ii) both $i_{1}$ and $i_{2}$ are matched to $s_{2}$. In the first case, let without loss of generality $i_{1}$ be matched to $s_{2}$ and $i_{2}$ be matched to $s_{1}$ and. Then, if $i_{1}$ prefers $s_{1}$ and $i_{2}$ prefers $s_{2}$, swapping their assignments would Pareto dominate the outcome produced by the mechanism while still respecting modular upper-bounds.
@@
In every case, the following occurs: either the outcome produced by the mechanism (which is constrained to be upper–bound respecting) violates the modular upper–bound (by assigning more than one officer to $s_{1}$), or there exists another allocation that both respects the upper–bounds and yields a strict Pareto improvement with respect to the officers’ true preferences. Therefore, for these modular upper-bounds, no mechanism exists that is simultaneously visibly fair and is constrained Pareto efficient.
@@ A.7 Proof of Theorem 8
We show by induction on the assignment order that for any modular upper-bound $(\Xi_{h},S_{h},k_{h})$, the number of officers with type in $\Xi_{t}$ assigned to states in $S_{h}$ never exceeds $k_{h}$.
Base Case. Before any assignment (step 0), no officer is assigned and the count is $0\leq k_{h}$.
Inductive Step. Suppose that after assigning officers $i_{1},\dots,i_{k-1}$, the upper-bound $(\Xi_{h},S_{h},k_{h})$ is not violated. When officer $i_{k}$ (of type $t_{i_{k}}$) is considered, if $t_{i_{k}}\not\in\Xi_{h}$ the claim is unaffected. If $t_{i_{k}}\in\Xi_{h}$ and $i_{k}$ is assigned a state $s$ with $s\in S_{h}$, then by construction $s$ belongs to zone $z_{1}$ for $i_{k}$ and the current count of officers with type $t\in\Xi_{h}$ in $S_{h}$ is strictly lower than $k_{h}$. If after assignment the count reaches $k_{h}$, the mechanism updates $B_{\cdot}^{t}$ by adding $h$; hence, any subsequent officer of type $t$ will have $S_{h}$ excluded from her zone $z_{1}$. Thus, by induction, no upper-bound is ever violated.
@@
    Binding Upper-Bound: The other reason for $a^{\prime}_{i^{*}}$ not to be in $i^{*}$’s zone $z_{1}$ under $a$ is if $a^{\prime}_{i^{*}}$ belonged to a set $S_{h}$ for which the corresponding upper-bound $(\Xi_{h},S_{h},k_{h})$ was already binding at the time of $i^{*}$’s assignment. However, if $a^{\prime}_{i^{*}}$ were in such a set and $a^{\prime}$ respects the upper-bound, then $a^{\prime}_{i^{*}}$ could not be assigned to $i^{*}$ without causing the total number of type $t_{i^{*}}$ officers in $S_{h}$ to exceed $k_{h}$.